# Multi-Scale Fusion for Object Representation

### v1

@@ Abstract
Representing images or videos as object-level feature vectors, rather than pixel-level feature maps, facilitates advanced visual tasks. Object-Centric Learning (OCL) primarily achieves this by reconstructing the input under the guidance of Variational Autoencoder (VAE) intermediate representation to drive so-called slots to aggregate as much object information as possible. However, existing VAE guidance does not explicitly address that objects can vary in pixel sizes while models typically excel at specific pattern scales. We propose Multi-Scale Fusion (MSF) to enhance VAE guidance for OCL training. To ensure objects of all sizes fall within VAE’s comfort zone, we adopt the image pyramid, which produces intermediate representations at multiple scales; To foster scale-invariance/variance in object super-pixels, we devise inter/intra-scale fusion, which augments low-quality object super-pixels of one scale with corresponding high-quality super-pixels from another scale. On standard OCL benchmarks, our technique improves mainstream methods, including state-of-the-art diffusion-based ones. The source code is available in the supplemental material.
@@ 1 Introduction
Object-level representation learning is highly influenced by textures. Early OCL methods, such as the mixture-based (Locatello et al., 2020; Kipf et al., 2022; Elsayed et al., 2022), which directly reconstruct the input pixels as supervision, often struggle with objects that have complex textures. To address this issue, mainstream OCL methods, such as transformer-based (Singh et al., 2022a; c), foundation-based (Seitzer et al., 2023; Zadaianchuk et al., 2024) and diffusion-based (Jiang et al., 2023; Wu et al., 2023), leverage well-learnt intermediate representation for guided reconstruction, so as to drive slots to aggregate as much object-level information as possible.
Objects can appear at different scales in images or videos due to variations in imaging distance and their actual sizes. At the same time, encoders/decoders usually perform optimally at specific pattern scales, thus ”divide-and-rule” is necessary like in object detection and instance/panoptic segmentation (Zou et al., 2023; Minaee et al., 2021). Mainstream OCL methods, however, do not explicitly provide such multi-scale guidance to object representation. This is a long-overlooked bottleneck that limits mainstream OCL performance.
@@
Figure 1: Multi-scale fusion is necessary in VAE guidance to OCL training. Upper right: simplified architecture of OCL with VAE guidance. Upper left: single-scale VAE encoding/decoding only represent patterns of some certain scale in high quality/confidence. Lower left and right: by resizing the input into different sizes, patterns of different scales all fall within the VAE encoder/decoder’s comfort-zone. For instance, in scale 1, the low quality super-pixels of objects 2 and 3 can be augmented by their corresponding high quality super-pixels in scales 2 and 3, respectively.
@@
Suppose there are three objects of different sizes in the input image or video frame. As shown in Fig. 1 upper left, current single-scale VAE encoding/decoding are likely only good at some specific scale of patterns, and only represent object 1 in high quality or confidence, leaving the other two being represented in low quality/confidence. Thus the intermediate representation of such single-scale VAE could not provide good guidance to OCL training on objects 2 and 3. But, as shown in Fig. 1 lower left and right, if we resize the input into three scales, we can ensure all the objects fall within the comfort-zone of the VAE encoder/decoder. Moreover, as shown in Fig. 1 lower left center, for scale 1, the low-quality super-pixels of object 2 and 3 can be augmented by their corresponding representations in scale 2 and 3, respectively. This is our motivation.
@@
We propose Multi-Scale Fusion (MSF) upon VAE intermediate representation, to guide OCL of both the transformer-based classics and diffusion-based state-of-the-arts. (i) We adopt the image pyramid technique, which resizes input images or video frames into different scales, to ensure objects of various sizes always fall within VAE encoder/decoder’s comfort zones. (ii) We devise inter/intra-scale quantized fusion, which quantize VAE intermediate representation of different scales with shared/specified codebooks, to foster scale-invariance/variance in object representation. (iii) Our technique augments VAE intermediate discrete representation with better object separability and, in turn, provides better guidance for OCL training.
@@ 2 Related Work
VAE & OCL. Pretrained VAE models (Im Im et al., 2017; Van Den Oord et al., 2017), like dVAE or VQ-VAE, are often used to provide well-learnt intermediate representations that highlight object separability in the input’s feature maps by suppressing texture redundancies for OCL. With such guidance, transformer-based OCL methods (Singh et al., 2022a; c), like SLATE and STEVE, decode slots, which are extracted by SlotAttention (Locatello et al., 2020) from input images or video frames, into the input via a transformer decoder. Similarly, diffusion-based methods (Jiang et al., 2023; Wu et al., 2023), like SlotDiffusion, decode slots via a diffusion model. These architectures drive slots to aggregate as much object information as possible, forming good object representation. Our work is conducted upon these two lines of work. And foundation-based methods (Seitzer et al., 2023; Zadaianchuk et al., 2024), like DINOSAUR, are also included for complete comparison.
Channel Grouping. Grouping features along the channel dimension helps to learn expressive representations, making it widely adopted in CNNs (Zhang et al., 2018) and ViTs (Gu et al., 2022). Grouping can be performed directly on feature map channels (Krizhevsky et al., 2012; Huang et al., 2018; Chen et al., 2019), or indirectly on convolution weights (Zhao et al., 2021; 2022). This is also explored in OCL setting. SysBinder (Singh et al., 2022b) groups slots’ queries to aggregate different object information, achieving emerging attributes under extra supervision. The most recent GDR (Zhao et al., 2024a) and OGDR (Zhao et al., 2024b) group the VAE codebook from feature level into combinable attribute level, achieving better generalizability and convergence in object representing. Our work also groups it but for scale-invariance and -variance.
Multi-Scale. Objects can exist at multiple scales due to variations in type and distance. Visual pattern processors like CNN (Zhang et al., 2018) and ViT (Gu et al., 2022) typically have their comfort zones as the scale distribution of objects in training data is always nonuniform. In computer vision tasks, like object detection (Zou et al., 2023) and instance/panoptic segmentation (Minaee et al., 2021), the multi-scale issue is significant, leading to the wide adoption of “divide-and-rule”. Multi-scale training using the image pyramid (Adelson et al., 1984) makes object scale distribution more uniform. Multi-scale designs like feature pyramid network (Lin et al., 2017) augment feature representations of different scales with one another, which are usually heavy and implicit. We address the multi-scale issue explicitly in the OCL setting for the first time.
@@ 3 Proposed Method
We propose Multi-Scale Fusion (MSF), to explicitly address the multi-scale issue in VAE intermediate representation, and thereby to guide Object-Centric Learning (OCL) better. Our technique is applicable to both transformer-based OCL (Singh et al., 2022a; c) classics and diffusion-based OCL (Wu et al., 2023) state-of-the-arts, which are our basis methods. Note that two types of methods use dVAE (Im Im et al., 2017) and VQ-VAE (Van Den Oord et al., 2017) respectively, but we unify them with VQ-VAE like that in OGDR (Zhao et al., 2024b).
@@ 3.2 Naive Multi-Scale Training
In images or videos, there can be objects of different pattern sizes due to their original sizes and distances to the imaging point; On the other hand, however, models processing them usually have their best-at scales of visual patterns because of the non-uniform object size distribution of their training data. An effective solution is image pyramid (Adelson et al., 1984), which resizes the input into different sizes for parallel processing via the model. During training, these naive multi-scale representations make the pattern scale distribution more uniform; during testing, this ensures different size objects fall within the model’s comfort zone.
Assume the input image or video frame is $\bm{T}_{1}:=\bm{T}$, and is down-sampled with typical bi-linear interpolation into $\bm{T}_{n}\in\mathbb{R}^{\frac{h_{0}}{2^{n-1}}\times\frac{w_{0}}{2^{n-1}}\times c_{0}}$, where $n=1...N$ means different spatial scales and we use $N=3$.
@@
Multi-scale VAE quantization is $f_{\mathrm{q}}=f_{\mathrm{m}}\circ f_{\mathrm{s}}:(\bm{Z}_{n},\bm{C})\rightarrow(\bm{P}_{n},\bm{I}_{n},\bm{X}_{n})$, where $\bm{P}_{n}\in\mathbb{R}^{\frac{h}{2^{n-1}}\times\frac{w}{2^{n-1}}\times m}$, $\bm{I}_{n}\in\mathbb{R}^{\frac{h}{2^{n-1}}\times\frac{w}{2^{n-1}}}$ and $\bm{X}_{n}\in\mathbb{R}^{\frac{h}{2^{n-1}}\times\frac{w}{2^{n-1}}\times c}$. Here codebook $\bm{C}$ is shared over all $N$ scales of representation so as to facilitate scale-invariant pattern learning that differentiates among different object types. But this strong inductive bias can lose scale-variant information that differentiates among different object instances of the same type; Besides, no information exchange among the scales is a kind of information waste – The object pattern that is well represented in one scale should augment its poorly represented patterns in the other scales. We will address this in the next subsection.
@@
During VAE pretraining, there are $N$ sets of reconstruction and quantization losses, each of which is defined identically to those in the former subsection. During OCL training, only scale $n=1$ is used as guidance, thus the supervision is no different from that in the former section. Although multi-scale guidance can be used, it is always harmful in practice. This can be attributed to that scales $n>1$ have too large down sampling and lose too much information to guide object-level information aggregation.
By following such recipe, hopefully at least one of the $N$ scales of representation provides high-quality representation to an arbitrary size object.
@@ 3.3 Inter/Intra-Scale Quantized Fusion
Based on the naive multi-scale quantization, we further devise inter/intra-scale fusion among those multi-scale representations. Inter-scale fusion utilizes scale-invariance to augment poorly represented features in some scales with the well-represented features in another scale; Intra-scale fusion augments the aforementioned scale-invariant features with the scale-variant features in the same scale. We refer to scale-invariance as patterns that differentiate among different object types, and scale-variance as patterns that distinguish among different object instances of the same type.
Firstly, we project VAE continuous intermediate representation of a certain scale $\bm{Z}_{n}$ into two copies, $\bm{Z}_{n}^{\mathrm{si}}$ for scale-invariance and $\bm{Z}_{n}^{\mathrm{sv}}$ for scale-variance, respectively. These two copies will ultimately be combined back after quantized fusion, thus we employ the invertible project-up-project-down design proposed in Zhao et al. (2024b) to maintain the inductive biases added in between.
@@
Secondly, for the scale-variant continuous representation, we conduct quantization by a codebook specific to the current scale. $f_{\mathrm{q}}=f_{\mathrm{m}}\circ f_{\mathrm{s}}:(\bm{Z}_{n}^{\mathrm{sv}},\bm{C}_{n}^{\mathrm{sv}})\rightarrow(\bm{P}_{n}^{\mathrm{sv}},\bm{I}_{n}^{\mathrm{sv}},\bm{X}_{n}^{\mathrm{sv}})$. These non-sharing codebooks allow for scale-variant representation learning.
Meanwhile, for the scale-invariant continuous representation, we conduct inter-scale fusion over scale-invariant quantized representations of all scales. $f_{\mathrm{rf}}:\{(\bm{Z}_{n}^{\mathrm{si}},\bm{C}^{\mathrm{si}})\}\rightarrow\{(\dot{\bm{P}}_{n}^{\mathrm{si}},\dot{\bm{I}}_{n}^{\mathrm{si}},\dot{\bm{X}}_{n}^{\mathrm{si}})\}$, where $n=1...N$. The shared codebook enforces scale-invariant representation learning.
Now that all scales $\{\bm{X}_{n}^{\mathrm{si}}\}$ consist of codes in a shared codebook $\bm{C}^{\mathrm{si}}$ with corresponding matching probabilities $\{\bm{P}_{n}^{\mathrm{si}}\}$, for the same (super-)pixel region under $N$ scales, the matching probabilities indicate how confident the model is about the representation quality. We utilize this to augment the low-confidence/quality representation in some scales with the high-confidence/quality representation in another scale, as below. Take scales $n-1$, $n$ and $n+1$ as an example:
@@
Lastly, for each scale $n$, conduct the intra-scale fusion between the fused scale-invariant representation and scale-variant representation. $f_{\mathrm{af}}:(\dot{\bm{X}}_{n}^{\mathrm{si}},\bm{X}_{n}^{\mathrm{sv}})\rightarrow\bm{X}_{n}$. This combines the augmented scale-invariant and scale-variant copies back into the final discrete intermediate representation. Note that this projection follow the aforementioned invertible project-up-project-down design.
@@ 3.4 Resource Consumption Analysis
Here we list our hyper-parameters and estimate the computation and memory consumption. We focus on the VAE part as our modifications to the basis methods only affect this part.
@@
However, for OCL training, heavy modules, like the primary encoder, SlotAttention and primary decoder, are added. Thus the parameter decrease and resource consumption inscrease are all negligible. Besides, no hyper-parameters in our technique requires extra tuning.
@@ 4 Experiments
4.1 OCL Performance
To evaluate the quality of the learnt object representation, we measure unsupervised segmentation accuracy using OCL’s byproduct segmentation masks. Like others, the metrics are ARI (Adjusted Rand Index)  [^1] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.adjusted_rand_score.html, ARIfg (foreground), mIoU (mean Intersection-over-Union)  [^2] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.jaccard_score.html and mBO (mean Best Overlap) (Caron et al., 2021). The datasets being used are either synthetic images, i.e., ClevrTex  [^3] https://www.robots.ox.ac.uk/~vgg/data/clevrtex/; or real-world images, i.e., COCO  [^4] https://cocodataset.org/#panoptic-2020 and VOC  [^5] http://host.robots.ox.ac.uk/pascal/VOC/voc2012/index.html; or synthetic videos, i.e., MOVi-C/D/E  [^6] https://github.com/google-research/kubric/tree/main/challenges/movi.
To evaluate how much performance our MSF can boost existing OCL methods, we use SLATE (Singh et al., 2022a) and STEVE (Singh et al., 2022c) to represent transformer-based classics, for images and videos respectively; use SlotDiffusion (Wu et al., 2023) to represent diffusion-based state-of-the-arts for images. We also include DINOSAUR (Seitzer et al., 2023) as a foundation-based representative. Similar improvers SysBinder (Singh et al., 2022b) and OGDR (Zhao et al., 2024b) are compared here too. We use DINO (Caron et al., 2021) as the primary encoder to make sure MSF is effective on very strong baselines.
Details about dataset processing, model architectures and training schemes are in the appendix.
As shown in Tab. 2, 2 and 4, MSF is beneficial to both transformer-based classics and diffusion-based state-of-the-arts. On complex synthetic dataset ClevrTex, MSF improves SLATE significantly, up to 12% in ARIfg. On challenging real-world dataset COCO, MSF still manages to improve SLATE by 6% in ARIfg. MSF even boosts SlotDiffusion, which is state-of-the-art, by 3% in ARI on VOC. Compared with other improvers, i.e., SysBinder@$g$2 and OGDR@$g$2, our MSF also beats or at least ties them on SLATE, while beats them consistently on STEVE. On SlotDiffusion, our MSF beats the strong competitor OGDR@$g$2, and defeats DINOSUAR.
@@ 4.1 OCL Performance
Table 4: Quantitative object separability of VAE guidance. “Inter-Cluster” is the mean distance of super-pixels within a cluster, while “Intra-Cluster” is that among different clusters. “@c256” and “@c4” are VAE intermediate channel dimensions, for SLATE/STEVE and SlotDiffusion respectively.
Therefore, our MSF is a general and effective improver to both transformer-based and diffusion-based OCL methods.

### v2

@@ Abstract
Representing images or videos as object-level feature vectors, rather than pixel-level feature maps, facilitates advanced visual tasks. Object-Centric Learning (OCL) primarily achieves this by reconstructing the input under the guidance of Variational Autoencoder (VAE) intermediate representation to drive so-called slots to aggregate as much object information as possible. However, existing VAE guidance does not explicitly address that objects can vary in pixel sizes while models typically excel at specific pattern scales. We propose Multi-Scale Fusion (MSF) to enhance VAE guidance for OCL training. To ensure objects of all sizes fall within VAE’s comfort zone, we adopt the image pyramid, which produces intermediate representations at multiple scales; To foster scale-invariance/variance in object super-pixels, we devise inter/intra-scale fusion, which augments low-quality object super-pixels of one scale with corresponding high-quality super-pixels from another scale. On standard OCL benchmarks, our technique improves mainstream methods, including state-of-the-art diffusion-based ones. The source code is available on https://github.com/Genera1Z/MultiScaleFusion.
@@ 1 Introduction
Object-level representation learning is highly influenced by textures. Early OCL methods, such as the mixture-based (Locatello et al., 2020; Kipf et al., 2022; Elsayed et al., 2022), which directly reconstruct the input pixels as supervision, often struggle with objects that have complex textures. To address this issue, mainstream OCL methods, such as transformer-based (Singh et al., 2022a; c), foundation-based (Seitzer et al., 2023; Zadaianchuk et al., 2024) and diffusion-based (Jiang et al., 2023; Wu et al., 2023b), leverage well-learnt intermediate representation for guided reconstruction, so as to drive slots to aggregate as much object-level information as possible.
@@
Figure 1: Multi-scale fusion is necessary in VAE guidance to OCL training. Upper right: simplified architecture of OCL with VAE guidance. Upper left: single-scale VAE encoding/decoding only represent patterns of some certain scale in high quality/confidence. Lower left and right: by resizing the input, patterns of different scales all fall within the VAE encoder/decoder’s comfort-zone; low-quality super-pixels in one scale can be augmented by the high-quality in other scales.
Objects appear at different scales in images or videos due to imaging distances and actual sizes. Encoders/decoders usually perform well on certain pattern scales, necessitating ”divide-and-rule” (Zou et al., 2023; Minaee et al., 2021). Mainstream OCL, however, overlooks such an issue.
@@
Suppose three objects of different sizes in the input image or video frame. As shown in Fig. 1 upper left, current single-scale VAE encoding/decoding are likely only good at some specific scale of patterns, and only represent object 1 in high quality or confidence, leaving the other two being represented in low quality. Thus such single-scale VAE representation could not provide good guidance to OCL training on objects 2 and 3. But, as shown in Fig. 1 lower left and right, if resize the input into three scales, we ensure all the objects fall within the comfort-zone of the VAE encoder/decoder. Moreover, as shown in Fig. 1 lower left center, for scale 1, the low-quality super-pixels of object 2 and 3 can be augmented by their corresponding representations in scale 2 and 3, respectively.
@@
We propose Multi-Scale Fusion (MSF) upon VAE representation, to guide OCL of both the transformer-based classics and diffusion-based state-of-the-arts. (i) We utilize technique image pyramid by resizing the input into different scales, to ensure objects of various sizes all fall within VAE encoder/decoder’s comfort zones. (ii) We devise inter-/intra-scale quantized fusion by quantizing VAE representation of different scales with shared/specified codebooks, to foster scale-invariance/-variance in object representation. (iii) Our technique augments VAE discrete representation with better object separability and, in turn, provides better guidance for OCL training.
@@ 2 Related Work
VAE & OCL. Pretrained VAE (Im Im et al., 2017; Van Den Oord et al., 2017), like dVAE or VQ-VAE, can provide representations that highlight object separability in the input by suppressing texture redundancies for OCL. With such guidance, transformer-based OCL methods (Singh et al., 2022a; c), like SLATE and STEVE, decode slots, which are extracted by SlotAttention (Locatello et al., 2020) from input images or video frames, into the input via a transformer decoder. Similarly, diffusion-based methods (Jiang et al., 2023; Wu et al., 2023b), like SlotDiffusion, decode slots via a diffusion model. Slots are driven to aggregate as much object information for good object representation. We improve these two lines of work. Foundation-based methods (Seitzer et al., 2023; Zadaianchuk et al., 2024), like DINOSAUR, are also included for complete comparison.
Channel Grouping. Widely adopted in CNNs (Zhang et al., 2018) and ViTs (Gu et al., 2022), grouping features along the channel dimension helps learn expressive representations. Grouping can be performed on features (Krizhevsky et al., 2012; Huang et al., 2018; Chen et al., 2019), or on weights (Zhao et al., 2021; 2022). SysBinder (Singh et al., 2022b) explores this in OCL by grouping slots queries to aggregate different object information, for emerging attributes under extra supervision. The most recent GDR (Zhao et al., 2024) groups the VAE codebook from features into combinatorial attributes, achieving better generalization and convergence in object representing. Our work also groups it but for scale-invariance and -variance.
Multi-Scale. Objects exist at different scales due actual sizes and imaging distances. CNN (Zhang et al., 2018) and ViT (Gu et al., 2022) typically have their comfort zones as the scale distribution of objects in training data is always nonuniform. In computer vision, like detection (Zou et al., 2023) and segmentation (Minaee et al., 2021), the multi-scale issue is significant, leading to the wide adoption of “divide-and-rule”. The image pyramid (Adelson et al., 1984) produces multi-scale representations by resizing; Feature pyramid network (Lin et al., 2017) and variants mix features of different layers by channel-concat and element-wise-sum. We are the first to realize multi-scale fusion on VAE, augmenting every scale with the high-quality representation in other scales.
@@ 3 Proposed Method
We propose Multi-Scale Fusion (MSF), to explicitly address the multi-scale issue in VAE intermediate representation, and thereby to guide Object-Centric Learning (OCL) better. Our technique is applicable to both transformer-based OCL (Singh et al., 2022a; c) classics and diffusion-based OCL (Wu et al., 2023b) state-of-the-arts, which are our basis methods. Note that two types of methods use dVAE (Im Im et al., 2017) and VQ-VAE (Van Den Oord et al., 2017) respectively, but we unify them with VQ-VAE like that in GDR (Zhao et al., 2024).
@@ 3.2 Naive Multi-Scale Training
There can be objects of different sizes in images or videos, while models usually have best-at patterns scales due to non-uniform data distributions. This makes image pyramid (Adelson et al., 1984) an effective technique, where the input is resized into different sizes and processed in parallel. Such multi-scale representations make the pattern scale distribution more uniform for training, and ensures different sized objects to fall within models’ comfort zone for testing.
Input image or video frame $\bm{T}_{1}:=\bm{T}$ is down-sampled with typical bi-linear interpolation into $\bm{T}_{n}\in\mathbb{R}^{\frac{h_{0}}{2^{n-1}}\times\frac{w_{0}}{2^{n-1}}\times c_{0}}$, where $n=1...N$ means different spatial scales and we use $N=3$.
@@
Multi-scale VAE quantization is $f_{\mathrm{q}}=f_{\mathrm{m}}\circ f_{\mathrm{s}}:(\bm{Z}_{n},\bm{C})\rightarrow(\bm{P}_{n},\bm{I}_{n},\bm{X}_{n})$, where $\bm{P}_{n}\in\mathbb{R}^{\frac{h}{2^{n-1}}\times\frac{w}{2^{n-1}}\times m}$, $\bm{I}_{n}\in\mathbb{R}^{\frac{h}{2^{n-1}}\times\frac{w}{2^{n-1}}}$ and $\bm{X}_{n}\in\mathbb{R}^{\frac{h}{2^{n-1}}\times\frac{w}{2^{n-1}}\times c}$. Here codebook $\bm{C}$ is shared over all $N$ scales so as to facilitate scale-invariant pattern learning that differentiates among different object types. But this strong inductive bias can lose scale-variant information that differentiates among different object instances of the same type; Besides, no information exchange among the scales is a kind of information waste – The object pattern that is well represented in one scale should augment its poorly represented patterns in the other scales. We address this in the next subsection.
@@
During VAE pretraining, there are $N$ sets of reconstruction and quantization losses. During OCL training, only scale $n=1$ is used as guidance, thus the supervision is no different from the former subsection. Although multi-scale guidance can be used, it is always harmful in practice.
By following such recipe, hopefully at least one of the $N$ scales of representation provides high-quality representation to some sized objects.
@@ 3.3 Inter/Intra-Scale Quantized Fusion
Based on the above multi-scale quantization, we further devise our novel fusion among the scales. Inter-scale fusion utilizes scale-invariance to augment poorly represented features in some scales with the well-represented features in other scales; Intra-scale fusion supplements with scale-variant details in the corresponding scale. Scale-invariant patterns differentiate among different object types, while scale-variance distinguishes among different object instances of the same type.
Firstly, we project VAE continuous representation of a certain scale $\bm{Z}_{n}$ into two copies, $\bm{Z}_{n}^{\mathrm{si}}$ for scale-invariance and $\bm{Z}_{n}^{\mathrm{sv}}$ for scale-variance, respectively. These two will finally be combined back, thus we employ the invertible project-up-project-down design proposed in Zhao et al. (2024) to maintain the inductive biases added in between.
@@
Secondly, for the scale-variant continuous representation, we conduct quantization by a codebook specific to the current scale. $f_{\mathrm{q}}=f_{\mathrm{m}}\circ f_{\mathrm{s}}:(\bm{Z}_{n}^{\mathrm{sv}},\bm{C}_{n}^{\mathrm{sv}})\rightarrow(\bm{P}_{n}^{\mathrm{sv}},\bm{I}_{n}^{\mathrm{sv}},\bm{X}_{n}^{\mathrm{sv}})$. These non-sharing codebooks enforces scale-variant representation learning.
Meanwhile, for the scale-invariant continuous representation, we conduct inter-scale fusion over scale-invariant quantization of all scales. $f_{\mathrm{rf}}:\{(\bm{Z}_{n}^{\mathrm{si}},\bm{C}^{\mathrm{si}})\}\rightarrow\{(\dot{\bm{P}}_{n}^{\mathrm{si}},\dot{\bm{I}}_{n}^{\mathrm{si}},\dot{\bm{X}}_{n}^{\mathrm{si}})\}$, where $n=1...N$. The shared codebook enforces scale-invariant representation learning.
Now that all scales $\{\bm{X}_{n}^{\mathrm{si}}\}$ consist of codes in a shared codebook $\bm{C}^{\mathrm{si}}$, matching probabilities $\{\bm{P}_{n}^{\mathrm{si}}\}$ for the same (super-)pixel region under different scales indicate how confident the model is about the representation quality. We utilize this to augment the low-quality representation in some scales with the high-quality in other scales. Take scales $n-1$, $n$ and $n+1$ as an example:
@@
Lastly, for each scale $n$, conduct the intra-scale fusion between the fused scale-invariant representation and scale-variant representation. $f_{\mathrm{af}}:(\dot{\bm{X}}_{n}^{\mathrm{si}},\bm{X}_{n}^{\mathrm{sv}})\rightarrow\bm{X}_{n}$. This combines the augmented scale-invariant and scale-variant copies back into the final discrete intermediate representation.
@@ 4 Experiments
Table 4: Object separability of VAE guidance. “Inter-Cluster” is the mean distance of super-pixels within a cluster, while “Intra-Cluster” is that among different clusters. “@c256” and “@c4” are VAE intermediate dimensions, for SLATE/STEVE and SlotDiffusion, respectively.
4.1 OCL Performance
To evaluate the quality of OCL object representations, i.e., slots, we use the accuracy of OCL’s byproduct (unsupervised) segmentation as an intuitive measurement. Metrics including ARI (Adjusted Rand Index) [^1] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.adjusted_rand_score.html, ARIfg (foreground), mIoU (mean Intersection-over-Union) [^2] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.jaccard_score.html and mBO (mean Best Overlap) (Caron et al., 2021) are used to measure OCL’s byproduct segmentation accuracy. Datasets are either synthetic images, i.e., ClevrTex [^3] https://www.robots.ox.ac.uk/~vgg/data/clevrtex/; or real-world images, i.e., COCO [^4] https://cocodataset.org/#panoptic-2020 and VOC [^5] http://host.robots.ox.ac.uk/pascal/VOC/voc2012/index.html; or synthetic videos, i.e., MOVi-C/D/E [^6] https://github.com/google-research/kubric/tree/main/challenges/movi.
To evaluate how MSF boosts OCL, we use transformer-based classics, SLATE (Singh et al., 2022a) for images and STEVE (Singh et al., 2022c) for videos; use diffusion-based state-of-the-arts, SlotDiffusion (Wu et al., 2023b) for images. The foundation-based DINOSAUR (Seitzer et al., 2023) and competitors SysBinder (Singh et al., 2022b) and GDR (Zhao et al., 2024) are compared too. All primary encoders are unified as DINO (Caron et al., 2021) to form strong baselines.
Our MSF is a general improver to both transformer-based classics and diffusion-based state-of-the-arts, as shown in Tab. 2, 2 and 4. On complex synthetic dataset ClevrTex, MSF improves SLATE significantly, up to 12% in ARIfg. On challenging real-world dataset COCO, MSF still manages to improve SLATE by 6% in ARIfg. MSF even boosts SlotDiffusion, which is state-of-the-art, by 3% in ARI on VOC. Compared with other improvers, i.e., SysBinder@$g$2 and GDR@$g$2, our MSF also beats or at least ties them on SLATE, while beats them consistently on STEVE. On SlotDiffusion, our MSF beats the strong competitor GDR@$g$2, and defeats DINOSUAR.